Find the solution of the given differential equation and then find the particular solution for which a point (x,y) is given:


dy/dx = x


1
Expert's answer
2021-08-31T12:59:38-0400

"\dfrac{dy}{dx}=x"

"dy=xdx"

Integrate


"\int dy=\int xdx"

"y=\dfrac{1}{2}x^2+C"

A point "(x_0, y_0)" is given.


"y_0=\dfrac{1}{2}x_0^2+C=>C=y_0-\dfrac{1}{2}x_0^2"

 The particular solution is


"y=y_0+\dfrac{1}{2}x^2-\dfrac{1}{2}x_0^2"